# Enhanced Single-Photon Detector: Achieving Superconducting-Level Performance with Conventional Quantum Technology

### v1

@@ Abstract
High-performance single-photon detectors (SPDs) are indispensable components for quantum optical tasks. However, the reliance of state-of-the-art devices on superconducting materials imposes severe technological demands and challenging operational conditions (e.g., cryogenics), which hinder scalable commercial deployment. To address this, we propose the Enhanced Single-Photon Detector (ESPD) framework, a novel paradigm for achieving high-performance SPDs through the iterative enhancement of low-technology SPDs. Utilizing only standard quantum optical components, the ESPD scheme transforms a legacy non-superconducting SPD, with detection efficiency (DE) about $59\%$ and dark count rate (DCR) $10^{-2}$, into a device with superior performance metrics, achieving DE higher than $95\%$ and DCR below $10^{-9}$. This level of performance is comparable to or surpasses recently designed superconducting SPDs, allowing the minimal tolerable channel transmission rate for Quantum Key Distribution (QKD) protocols to be reduced by several orders of magnitude. Furthermore, the scheme’s device requirements are moderate, relying on readily available current technology, which ensures near-term experimental feasibility. The ESPD framework thus provides a clear, scalable path toward the large-scale deployment of high-performance SPDs and the commercialization of quantum communication technologies.
@@
Single photon-detector, Detective efficiency, Dark count rate, Quantum optics, Quantum key distribution
@@ 1 Introduction
Single-photon detectors (SPDs) play a pivotal role in quantum optics, as measurement is indispensable for information readout across virtually all applications. The two most critical performance metrics of an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$. The former influences the efficiency of nearly all quantum operations, while the latter constitutes the ultimate limitation in high-attenuation scenarios. Regarding DE, numerous quantum experiments rely on SPD signals to retrieve results. In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the success rate decreases exponentially with low-efficiency SPDs [1, 2, 3]. In Bell tests for quantum non-locality, verifying violations of the Clauser–Horne–Shimony–Holt or Eberhard inequalities routinely requires DE exceeding specific thresholds to close loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12]. In heralded single-photon sources, where entangled pairs are generated and one photon is detected to confirm the presence of the other, the efficiency of the SPD directly dictates the success rate [13, 14, 15, 16]. Regarding DCR, a primary constraint arises in quantum communication, particularly in quantum key distribution (QKD). The security of a practical QKD protocol relies on maintaining the quantum bit error rate (QBER) below a protocol-specific threshold, which is significantly lower than 50% [17, 18, 19, 20]. However, communication channels induce exponential photon attenuation [21, 22, 23, 24]. If the effective signal rate, determined by the channel transmission rate and thus communication distance, falls below the DCR of the SPDs, the QBER rises precipitously, as dark count events (which yield a 50% error rate) dominate the detection statistics. Consequently, the practical secure distance of QKD is strictly limited by the DCR of the SPDs [25, 26, 27].
@@
Given their critical importance, the development of SPDs has been a continuous research focus since the inception of quantum technologies. While capable of operating at room temperature, traditional SPDs often suffer from both low DE and high DCR. For instance, a frequency upconversion SPD presented in 2004 operates at 300 K with a DE of approximately 59% and a DCR of $10^{-2}$ [28], while an InGaAs/InP detector presented in 2017 operates at 223 K with a DE of 27.5% and a DCR of $10^{-6}$ [29]. These specifications are insufficient for advanced quantum tasks, which typically require the DE of SPD above 90% with DCR below $10^{-7}$. To suppress DCR and enhance DE, recent efforts have shifted toward designing SPDs using superconducting devices, such as superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These detectors achieve performance superior to traditional counterparts, boosting DE to over 90% while suppressing DCR by several orders of magnitude. This capability enables the implementation of demanding quantum tasks, such as long-distance QKD distance [19, 34, 35, 36]. A summary of recent representative advancements in superconducting SPDs is presented in Table 1.
@@
However, as illustrated in the table, the reliance on superconductors necessitates stringent operating conditions, specifically with extremely low temperatures. Furthermore, these systems often entail high technical complexity and cost, hindering scalable commercial deployment.
To address these limitations, we propose a design for SPDs that meets advanced performance requirements, and even potentially outperforms superconducting devices, while relying on accessible, non-superconducting technology suitable for room-temperature operation. The proposed method, termed the Enhanced SPD (ESPD) scheme, iteratively upgrades the performance of an existing SPD level by level. As a demonstration of its capability, the ESPD scheme can enhance a non-superconducting SPD designed 20 years ago [28] to one achieving a DE exceeding 95% and a DCR below $10^{-9}$, thereby rivaling state-of-the-art devices. Our method not only provides a high-performance SPD design realizable with current (even commercial) technologies but also establishes a general framework for engineering high-performance SPDs using low-technology components. This opens a new avenue for research into SPDs and related quantum optical technologies.
2 Results
In this section, we detail the architecture of the Enhanced Single-Photon Detector (ESPD) scheme, which iteratively improves performance starting from a base (existing) SPD. Then we establish the theoretical framework governing its operation and provide analytical approximations to characterize its performance.
@@ 2.1 Design of the Enhanced Single-Photon Detector
The ESPD architecture is designed to upgrade the capabilities of a given SPD level by level, ultimately achieving high DE and low DCR.
For clarity, we define the DE as the probability of generating a positive detection event given a single-photon input, and the DCR as the probability of a positive report given a vacuum input. Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with DE $\eta_{s}$ and DCR $d_{s}$. The iteration begins with an initial detector, $\text{ESPD}_{0}$, characterized by $\eta_{0}=\eta$ and $d_{0}=d$.
@@
    DOF Selection: Select a manipulable DOF of the incoming photon (e.g., polarization). Let $\{|0\rangle,|1\rangle\}$ be the computational basis for this DOF. We treat the signal as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle,|1\rangle$ denote the orthonormal states of a non-vacuum photon. Note that if the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents non-vacuum. However, for practical implementations, it is often more convenient to use an internal DOF like polarization distinct from it.
@@
    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detections are registered among the $n_{s+1}+1$ detectors (comprising $n_{s+1}$ auxiliary detectors and one signal path detector). Otherwise, a negative report is returned. Here, $k_{s+1}$ is a pre-selected threshold integer. Here, $n_{s+1}$ should refer to the number of effective operations that survive post-selection (if the controlled gates include such), namely, the decision logic is essentially that the positive report rate on the multiple detections reaches a pre-defined threshold.
2.2 Theoretical Formulation
@@ 2.2.1 Notations
We define the parameters governing the system dynamics. As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively. Although $\eta_{s}$ may theoretically depend on the photon number distribution, we approximate it as a unified value for the given system. The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled gates and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
We assume all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$ for convenience. Let $p$ be the transmission probability of the signal path through a single module (an optional state preparation followed by a $C(F)$). Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input, accounting for loss but irrespective of decoherence (since the state can be re-prepared). Furthermore, let $Q$ represent the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement, but not including detection). Given the maturity of commercial linear optical components, we assume $Q\ll 1$ (typically $Q<10^{-3}$). Finally, let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively. For simplicity, we approximately assume $p$, $P$, and $Q$ are invariant across all $C(F)$ operations.
@@ 2.2.2 Assumptions
In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$. Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds. We further assume $d_{s}\ll P\eta_{s}$. This condition is trivial, for if it were violated, the signal contribution from the controlled operations would be indistinguishable from device noise, rendering the operations practically insignificant. Additionally, the transmission rate is assumed to be $p\approx 1$ (see Section 3.1). Finally, considering practical implementation costs, we assume $n_{s}$ and $k_{s}$ are moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
@@ 2.2.3 Formulations
Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum or vacuum input to the $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector. These are expressed as:
@@
Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum or vacuum output from the final gate, respectively, then:
@@
where the first and second terms correspond to scenarios where the signal path detector reports negative or positive, respectively.
If the photon survives all $C(F)$ operations, the probability is:
@@
where the mapping $G$ is defined by Eq. (1) to (7).
2.3 Analytical Approximation and Stability Analysis
@@ 2.4 Performance Simulations
While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD is mathematically complex, numerical methods provide a robust and sufficient approach for practical design. In this section, we present numerical simulations to validate the ESPD paradigm, demonstrating that substantial performance gains are achievable even with moderate initial hardware specifications. For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $95\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
Simulation Parameters: We evaluated the ESPD scheme using two baseline detectors: one representing older technology with $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], and another representing typical InGaAs/InP performance with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29]. The transmission probability is set to $p=0.99$, and the auxiliary path error rate $Q=0.2\%$ (see Section 3.1). We considered two controlled gate fidelity scenarios: $P=0.81$, representing standard linear optical C-NOT gates [38], and P=0.95P=0.95111Corresponding to about 0.22 dB, which is very likely to be achieved commercially in the short term, since it is confident to achieve by the commercial platform Q.ANT in 2026., representing high-fidelity implementations [39, 40]. Also, a further implementation with $P$ reduced to 0.4 is reported in Section 3.1. Various combinations of the configuration parameters $n_{s}$ and $k_{s}$ are evaluated.
Performance Outcomes: Table 2 to 5 and Fig. 2 to 5 summarize the achievable performance metrics across varying configurations and baseline architectures. The results indicate that the ESPD scheme significantly enhances the capabilities of conventional SPDs. Notably, with only three levels of enhancement, the detectors achieve performance metrics (DE $95\%$, DCR $10^{-9}$) that rival or surpass those of advanced superconducting SPDs.
3 Discussion
We now discuss the device requirements, potential applications, as well as current limitations and avenues for future work.
Table 2: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and high controlled gate fidelity, $P=0.99$.
@@ 3 Discussion
1 (4,2) $(88.7\%,\ 1.2\times 10^{-3})$ (6,3) $(86.2\%,\ 4.5\times 10^{-5})$ (7,3) $(90.9\%,\ 7.2\times 10^{-5})$
2 (4,2) $(98.6\%,\ 6.5\times 10^{-5})$ (6,3) $(97.4\%,\ 1.1\times 10^{-8})$ (7,3) $(97.7\%,\ 2.4\times 10^{-7})$
3 (4,2) $(99.0\%,\ 2.5\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
4 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
5 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
6 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
7 (4,2) $(99.0\%\ ,2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
8 (4,2) $(99.0\%,\ 2.4\times 10^{-5})$ (6,3) (97.9%, $9\%,\ 1.5\times 10^{-7}) (7$,3) $(97.9\%,\ 2.6\times 10^{-7})$
Para 4 Para 5 Para 6
[…]
1 (8,4) $(84.3\%,\ 1.8\times 10^{-6})$ (8,2) $(97.3\%,\ 4.2\times 10^{-3})$ (8,2) $(97.3\%,\ 4.2\times 10^{-3})$
2 (8,4) $(96.2\%,5.7\times 10^{-10})$ (8,4) $(96.9\%,\ 1.5\times 10^{-7})$ (8,3) $(97.9\%,\ 1.7\times 10^{-5})$
3 (8,4) $(96.8\%,9.5\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,\ 1.1\times 10^{-9})$
4 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
5 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
6 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
7 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
8 (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$ (8,4) $(96.9\%,9.8\times 10^{-10})$
(a) Detection efficiency $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) Dark count rate $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 2: Performance scaling for initial parameters $\eta_{0}=59.0\%$, $d_{0}=1.0\times 10^{-2}$, with $P=0.99$.
Table 3: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$, and moderate controlled gate fidelity, $P=0.81$.

### v2

# Enhanced Single-Photon Detector: A framework for Superconducting-Level Performance without cryogenic cooling

@@ Abstract
High-performance single-photon detectors (SPDs) are indispensable components for a wide range of quantum optical applications. However, the reliance of state-of-the-art devices on superconducting materials imposes severe technological demands and necessitates challenging operational conditions, such as cryogenics, thereby hindering scalable implementation. To address this, we propose the Enhanced Single-Photon Detector (ESPD) framework, a novel paradigm for achieving high-performance SPDs through the iterative enhancement of conventional room-temperature SPDs. Relying entirely on non-superconducting components and eliminating cryogenic cooling requirements, the ESPD scheme can upgrade a legacy SPD, with detection efficiency (DE) about $59\%$ and dark count rate (DCR) of $10^{-2}$, to a device with superior performance metrics, achieving DE exceeding $93\%$ and DCR below $10^{-9}$. This performance rivals or even surpasses that of state-of-the-art superconducting SPDs, allowing the minimal tolerable channel transmission rate for quantum key distribution (QKD) protocols to be reduced by several orders of magnitude. Although the architecture requires substantial integration effort, the scheme can provide superconducting-level performance without cryogenic cooling, offering a clear path toward the widespread deployment of high-performance SPDs as well as related quantum technologies in infrastructure-constrained environments.
@@
Room-Temperature Single photon-detector, Detective efficiency, Dark count rate, Quantum optics, Quantum key distribution
@@ 1 Introduction
Single-photon detectors (SPDs) play a pivotal role in quantum optics, as measurements are indispensable for information readout across virtually all applications. The two most critical performance metrics of an SPD are the detection efficiency (DE), denoted by $\eta$, and the dark count rate (DCR), denoted by $d$. The former affects the efficiency of nearly all quantum operations, while the latter imposes the ultimate limitation in high-attenuation scenarios. In terms of DE, numerous quantum experiments rely on SPD signals to retrieve results. In linear optical quantum computation, where success events are probabilistic and require the simultaneous detection of multiple photons, the success rate decreases exponentially with low-efficiency SPDs [1, 2, 3]. In Bell tests for quantum non-locality, verifying violations of the Clauser–Horne–Shimony–Holt or Eberhard inequalities routinely requires DE exceeding specific thresholds to close loopholes [4, 5, 6, 7, 8, 9, 10], such as approximately 67% [11] or 83% [12]. In heralded single-photon sources, where entangled pairs are generated, and one photon is detected to confirm the presence of the other, the efficiency of the SPD directly dictates the success rate [13, 14, 15, 16]. With respect to DCR, a primary constraint arises in quantum communication, particularly in quantum key distribution (QKD). The security of a practical QKD protocol relies on maintaining the quantum bit error rate (QBER) below a protocol-specific threshold, significantly lower than 50% [17, 18, 19, 20]. However, communication channels induce exponential photon attenuation [21, 22, 23, 24]. If the effective signal rate, determined by the channel transmission rate and thus communication distance, falls below the DCR of the SPDs, the QBER rises precipitously, as dark count events (which yield a 50% error rate) dominate the detection statistics. Consequently, the practical secure distance of QKD is strictly limited by the DCR of the SPDs [25, 26, 27].
@@
Driven by their critical importance, SPD has remained a continuous research focus since the inception of quantum technologies. While capable of operating at room temperature, traditional SPDs often suffer from both low DE and high DCR. For instance, a frequency upconversion SPD presented in 2004 operates at 300 K with DE about 59% and DCR of $10^{-2}$ [28], while an InGaAs/InP SPD presented in 2017 operates at 223 K with DE about 27.5% and DCR of $10^{-6}$ [29]. Such specifications are insufficient for advanced quantum tasks, which typically require the DE of SPD above 90% with DCR below $10^{-7}$. To suppress DCR and enhance DE, recent efforts have shifted toward designing SPDs using superconducting devices, such as superconducting nanowire SPDs (SNSPDs) [30, 31, 32, 33]. These SPDs achieve performance superior to traditional counterparts, boosting DE to over 90% while suppressing DCR by several orders of magnitude. This capability enables the implementation of demanding quantum tasks, such as long-distance QKD [19, 34, 35, 36]. A summary of recent representative advancements in superconducting SPDs is presented in Table 1.
@@
However, as illustrated in the table, the reliance on superconductors necessitates stringent operating conditions, specifically with extremely low temperatures. When these demanding conditions are relaxed, the SPD performance degrades significantly. This requirement introduces significant infrastructural requirements and operational complexity, fundamentally hindering the scalability and widespread implementation of high-performance quantum technology in field-deployable and environmentally constrained situations.
To overcome these limitations, we propose the enhanced SPD (ESPD) architecture that can meet advanced performance requirements, potentially outperforming superconducting ones, while relying solely on non-superconducting components suitable for room-temperature operation. The proposed method iteratively upgrades the performance of an existing conventional SPD level by level, ultimately to a high-performance one. As a demonstration of its capability, the ESPD scheme can enhance a non-superconducting SPD designed 20 years ago [28] with DE about 59% and DCR of $10^{-2}$ to one achieving DE exceeding 93% and DCR below $10^{-9}$, thereby can rival the state-of-the-art devices and reduce the minimal tolerable channel transmission rate in QKD by several orders of magnitude. Despite the substantial engineering integration effort required for its large-scale architecture, the ESPD method offers two key advancements: It not only provides a high-performance SPD design approach without cryogenic components, enabling a practical route toward accurate room-temperature photonic detection, but also establishes a paradigm shift for quantum information processing by leveraging multi-copy analysis and iterative enhancement to boost performance using conventional components. Consequently, this work opens a new avenue for the scalable deployment of high-performance SPDs and advances broader quantum optical technologies in infrastructure-constrained environments.
2 Result
In this section, we detail the architecture of the Enhanced Single-Photon Detector (ESPD) scheme, which iteratively improves performance initialized from a base (existing) SPD. We establish the theoretical framework governing its operation, provide analytical approximations to characterize its performance, and state its applications in QKD.
@@ 2.1 Design of the Enhanced Single-Photon Detector
The ESPD architecture is designed to recursively upgrade the capabilities of a given SPD level by level, ultimately achieving high DE and low DCR.
For clarity, we clarify that DE is the probability of generating a positive detection event given a non-vacuum (single-photon) input, and the DCR is the probability of a positive report given a vacuum input. Let $\text{ESPD}_{s}$ denote the ESPD at level $s$, with DE $\eta_{s}$ and DCR $d_{s}$. The iteration begins with an initial existing SPD, $\text{ESPD}_{0}$, characterized by initial value $\eta_{0}=\eta$ and $d_{0}=d$.
@@
    DOF Selection: Select a manipulable DOF of the incoming photon (e.g., polarization). Let $\{|0\rangle,|1\rangle\}$ denote the computational basis for this DOF. We treat the signal as a qutrit in the basis $\{|\emptyset\rangle,|0\rangle,|1\rangle\}$, where $|\emptyset\rangle$ denotes the vacuum state, and $|0\rangle,|1\rangle$ denote the orthonormal states of a non-vacuum photon [^1] Note that if the selected DOF is the photon number itself, $|0\rangle$ coincides with $|\emptyset\rangle$ and $|1\rangle$ represents non-vacuum, which degenerates to the qubit case. However, for practical implementations, it is often more convenient to use an internal DOF like polarization distinct from photon existence..
@@
    Decision Logic: A positive detection report for $\text{ESPD}_{s+1}$ is generated if and only if at least $k_{s+1}$ positive detections are registered among the $n_{s+1}+1$ detectors (comprising $n_{s+1}$ auxiliary detectors and one signal path detector). Otherwise, a negative report is returned. Here, $k_{s+1}$ is a pre-selected threshold integer, and $n_{s+1}$ refers to the number of effective operations [^2] For example, if the controlled gates are deterministic, then $n_{k+1}$ is exactly the number of used operations, while if the controlled gates are probabilistic, then $n_{k+1}$ refers to the successful ones after post-selection. Therefore, the decision logic is essentially that the positive report rate (among all effective reports) on the multiple detections reaches a pre-defined threshold..
2.2 Theoretical Formulations
@@ 2.2.1 Notations
We define the parameters governing the system dynamics. As stated, $\eta_{s}$ and $d_{s}$ represent the DE and DCR of $\text{ESPD}_{s}$, respectively. Although $\eta_{s}$ may depend on the photon number distribution, we approximate it as a unified value for the given system for simplicity. The parameters $n_{s+1}$ and $k_{s+1}$ denote the number of effective controlled operations and the decision threshold for $\text{ESPD}_{s+1}$, respectively.
For convenience, assume that all controlled operations $C(F_{i})$ are identical, denoted simply as $C(F)$. Let $p$ be the transmission probability of the signal path through a single module (an (optional) state preparation followed by a $C(F)$). Specifically, $p$ is the probability that the output signal remains non-vacuum given a non-vacuum input of a single non-cascaded module, accounting for loss but irrespective of decoherence (since the state can be re-prepared). Let $P$ and $Q$ denote the probabilities of obtaining a non-vacuum output on the controlled path before detection, given that the input to the corresponding module was non-vacuum or vacuum, respectively. Hence, $Q$ represents the intrinsic error rate of the auxiliary path in the absence of $C(F)$ (i.e., errors from auxiliary state preparation and measurement (SPAM), but not including detection). We approximately assume $p$, $P$, and $Q$ are invariant across all $C(F)$ module operations.
It is crucial to note that $p$, $P$, and $Q$ represent the intrinsic performance metrics of a single, non-cascaded controlled module. These parameters are defined at the component level and remain invariant with respect to the number of cascaded modules. The cumulative effects of loss and error due to the system’s cascaded architecture will be fully incorporated into the derived system equations, such as $\eta_{s+1}$ in Eq.(6).
@@ 2.2.2 Assumptions
In practical scenarios, SPD DE typically ranges from 10% to 90%, while the DCR is generally below $10^{-3}$. Consequently, the hierarchy $0\leq d\ll\eta\leq 1$ holds. We further assume $d_{s}\ll P\eta_{s}$, which is trivial: If it were violated, the signal contribution from the controlled operations would become indistinguishable from device noise, rendering the operations practically insignificant. Additionally, we assume that the transmission rate of a single controlled module $p\approx 1$, SPAM error of the auxiliary path $Q\ll 1$, given the maturity of commercial optical components (see Section 3.1). Moreover, for analytical simplicity, we assume the controlled gates are implemented in a feedforward controlled configuration. This ensures that $n_{s+1}$ represents both the number of effective modules and the total physical modules required for level $s+1$. However, we emphasize that the generalized ESPD design remains valid for post-selection controlled gates (for a brief discussion, see Section 3.1.2). Finally, considering practical implementation costs, we assume $n_{s}$ and $k_{s}$ are moderate integers satisfying $1\leq k_{s}\leq n_{s}$.
@@ 2.2.3 Formulations
Let $P_{s}$ and $Q_{s}$ denote the probabilities of obtaining a positive report on an auxiliary path given a non-vacuum or vacuum input to a $C(F)$ module, respectively, using $\text{ESPD}_{s}$ as the detector. They can be expressed as:
@@
Similarly, let $P^{\prime}_{s}$ and $Q^{\prime}_{s}$ denote the probabilities of obtaining a positive report on the signal path after all $C(F)$ operations, given a non-vacuum or vacuum output from the final controlled gate, respectively, then:
@@
where the first and second terms correspond to scenarios that the signal path detector reports negative and positive, respectively.
If the photon survives all $C(F)$ operations, the probability of a final positive report is:
@@
where the map $G_{s+1}$ is defined by Eq. (1) to (7).
2.3 Analytical Approximations and Stability Analysis
@@ 2.4 Performance Simulations
While obtaining an exact analytical solution for the non-linear, time-variant dynamic system governing the ESPD in Eq.(8) is mathematically complex, numerical methods provide a robust and sufficient approach for practical design. In this section, we present numerical simulations to validate the ESPD paradigm, demonstrating that substantial performance gains are achievable even with moderate initial hardware specifications. For example, a three-level ESPD can transform a semiconductor SPD into one exhibiting DE $93\%$ and DCR $10^{-9}$, effectively removing the necessity for superconducting materials while matching or exceeding the performance of state-of-the-art SNSPDs.
Simulation Parameters: We evaluated the performance of the ESPD scheme using two distinct baseline SPDs. The first represents an older technology, characterized by low performance parameters $(\eta_{0},d_{0})=(59\%,10^{-2})$ [28], while the second represents more modern InGaAs/InP SPD performance, with $(\eta_{0},d_{0})=(27.5\%,10^{-6})$ [29]. Additionally, the controlled-gate parameters are set to $p=0.98$, which can be achieved by commercial devices [38], and $P=0.97$, achievable recently by feedforward C-NOT gate implementations in photonic platforms [39, 40, 41, 42], while, for the auxiliary path, the error rate is set to $Q=0.2\%$, justified by the extremely low SPAM errors (without detections) achievable across specific DOF and the inherent design freedom in selecting the appropriate DOF for the ESPD implementation (see Section 3.1). We also implemented additional simulations using degraded parameter settings to evaluate robustness (also see Section 3.1).
Performance Outcomes: Achievable performance metrics across varying configurations and baseline architectures are summarized in Table 2 and 3 with visual representations in Fig. 2 and 3. The results indicate that the ESPD scheme can significantly enhance the capabilities of conventional SPDs. Notably, with only three or four levels of enhancement, the detectors achieve performance metrics (DE $93\%$, DCR $10^{-9}$) that rival or surpass those of advanced superconducting SPDs.
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 2: Performance for initial parameters $\eta_{0}=59.0\%$ and $d_{0}=1.0\times 10^{-2}$.
Table 2: Performance evolution of an ESPD initialized with a baseline detector from [28], with $\eta_{0}=59\%,\ d_{0}=10^{-2}$.
@@
1 (4,1) $(97.4\%,\ 5.3\times 10^{-2})$ (6,1) $(98.4\%,\ 7.5\times 10^{-2})$ (8,1) $(98.6\%,\ 9.5\times 10^{-2})$
2 (4,2) $(98.2\%,\ 2.7\times 10^{-2})$ (6,3) $(96.6\%,\ 1.2\times 10^{-2})$ (8,4) $(95.1\%,\ 7.4\times 10^{-3})$
3 (4,2) $(98.0\%,\ 7.7\times 10^{-3})$ (6,3) $(95.8\%,\ 8.9\times 10^{-5})$ (8,4) $(93.6\%,\ 8.1\times 10^{-7})$
4 (4,2) $(97.9\%,\ 8.4\times 10^{-4})$ (6,3) $(95.6\%,\ 1.7\times 10^{-7})$ (8,4) $(93.4\%,\ 8.6\times 10^{-10})$
5 (4,2) $(97.8\%,\ 5.6\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
6 (4,2) $(97.8\%,\ 2.5\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
7 (4,2) $(97.8\%\ ,2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
8 (4,2) $(97.8\%,\ 2.4\times 10^{-5})$ (6,3) $(95.6\%,\ 1.4\times 10^{-7})$ (8,4) $(93.4\%,\ 8.5\times 10^{-10})$
(a) DE $\eta_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
(b) DCR $d_{s}$ vs. level $s$ for varying $(n_{s},k_{s})$.
Figure 3: Performance for initial parameters $\eta_{0}=27.5\%$ and $d_{0}=1.0\times 10^{-6}$.
Table 3: Performance evolution of an ESPD initialized with a baseline detector from [29], with $\eta_{0}=27.5\%,\ d_{0}=10^{-6}$.